The rate of diffusion of a gas Z is twice (IV) oxide. [SO² =64]. What is the molecular mass of Z ?


1
Expert's answer
2021-03-01T03:25:06-0500

Rate of SO2= R

Rate of Z = 2R

Rate of Z/Rate of SO2 = √Molar mass of SO2/ √molar mass of Z


2R/R = √64/√Z

√Z = (R x √64)/ 2R


√Z = 4


Z= 4^2 = 16


Hence, the molecular mass of Z is 16g/mol